Sheryl says that the solutions of the quadratic equations w^2=49 and w^2+49=0 are the same. Do you agree with sheryl? Justify your answer.

Please heelp me

Answer by rothauserc(4718)

The solution for w^2 = 49 is
w = square root(49)
w = 7 or -7
now solution for w^2 + 49 = 0 is
subtract 49 from both sides of =
w^2 = -49
w = square root(-49)
w = 7 * square root(-1)
w = -7 * square root(-1)
now we write square root(-1) = i, therefore
w = 7i or -7i (these are imaginary numbers)
sheryl is not correct since we have different solutions for w^2=49 and w^2+49=0